OpenStudy (anonymous):

solve 3/11(w-2)-31/11=-3

OpenStudy (gw2011):

(3/11)(w-2)-31/11=-3 (3/11)w-6/11-31/11=-3 (3/11)w-37/11=-3 (3/11)w=-3+37/11 (3/11)w=-33/11+37/11 (3/11)w=4/11 w=(4/11)(113) w=4/3

OpenStudy (anonymous):

thank you

OpenStudy (gw2011):

You're welcome

OpenStudy (anonymous):

can you some how explain how you got the answer because my choices are 1 solution, 2a solution,or 0 sloution

OpenStudy (gw2011):

Multiply 3/11 times (w-2) to get (3/11)w-6/11. Now combine all like terms. You also have to find a common denominator for -3.

OpenStudy (gw2011):

There is one solution to the problem: w=4/3

OpenStudy (anonymous):

ok thank you so much

OpenStudy (gw2011):

You're welcome